an angle is 14 degree more than its complementary angles. what is its measure?

Found 2 solutions by ganeshindia, mananth:
Answer by ganeshindia(11)

Let the measure of the angle be x.
Therefore, its complement is 90 - x.
Given that,
angle = 14 degree + its complement
x = 14 + 90 -x
2x = 104
x = 104/2
x = 52 degrees.
therefore, the required angle is 52 degrees.

Answer by mananth(16949)

let angle be x
the complimentary = 90-x
x= (90-x)+14
x=104-x
add x to both sides
x+x=104-x+x
2x=104
divide by 2
2x/2 = 104/2
x=52 deg.